A rectangular field is 300 ft by 500 ft. A roadway of width x ft is to be built just inside the field. What is the widest the roadway can be and still leave 100,000 ft2 in the region?
Answer by richwmiller(17219)

I will assume that "in the region " means in the rectangular field.
Let's get the area of the field
300*500=150000
so the roadway shouldn't be more than 50000 ft^2
We aren't told the length nor whether the roadway goes around the perimeter or just across the field we don't know the length
Assuming the it is 300+500+500+300 long=1600 ft long
50000/1600= 31.25 ft wide